# Approximating the coefficients of the Bessel functions

### v2

@@ 1.1. Main result
Furthermore, Theorem˜1.1 generalizes the results of [limits_prob_measures] and resolves the question posed in the appendix of [matrix]. In particular, [limits_prob_measures] proves the implication of (b) from (a) after modifying the condition that $\lim_{N\rightarrow\infty}\frac{c_{\lambda}(N)}{(\theta N)^{\ell(\lambda)}}=0$ to $|c_{\lambda}(N)|=O(|\theta N|\max(N,|\theta N|)^{o_{N}(1)})$ for $\lambda\in\Gamma$ such that $\ell(\lambda)\geq 2$. The main contribution of this paper is the set of equivalence relations in the theorem where some conditions are based on scaling the coefficients of the logarithms of the input functions $F_{N}^{A}$, $F_{N}^{BC}$, and $F_{N}^{D}$ by varying powers of $\theta N$. The idea of scaling the coefficients by varying powers of $N$ while $\theta$ is fixed is mentioned in the appendix of [matrix]. However, in [limits_prob_measures], the coefficients are only scaled by $(\theta N)^{-1}$ and it is similarly the case that $\lim_{N\rightarrow\infty}|\theta N|=\infty$.
@@ 1.2. The $\theta N\rightarrow c\in\mathbb{C}$ regime
Note that [matrix] considers regime 3 for the root system $A^{N-1}$ while [rectangularmatrix] considers regime 4. The approach that we use to prove Theorem˜1.1 over regimes 1 and 2 is easily adaptable to proving similar results over regimes 3 and 4 as well. Theorems that we prove over regimes 3 and 4 are Theorem˜6.17, which generalizes the results of [matrix], and Theorem˜7.8, which generalizes the results of [rectangularmatrix].
@@ 1.4. Weak convergence to the free convolution
From [dunkloperators, dunklbessel], if $\theta\in\theta(\mathcal{R})$ such that $\theta(r)$ has nonnegative real part for all $r\in\mathcal{R}$, then $\theta\in\Theta(\mathcal{R})$, or equivalently, $\mathcal{D}(\mathcal{R}(\theta))$ is invertible. See Subsection˜2.1 for the definitions of these notions. The invertibility of the Dunkl operator implies the existence of it associated eigenfunctions, see Theorems˜2.5 and 2.8. In particular, this implies that the eigenfunctions exist in the context of Theorem˜1.2.
@@
Suppose $\geq 1, N\mathcal{R}\subset\mathbb{R}^{N}$ is a finite root system, and $\theta\in\theta(\mathcal{R})$ is nonnegative. Suppose $a_{1},a_{2}\in\mathbb{R}^{N}$. There exists a nonnegative probability measure $\mu^{\mathcal{R}(\theta)}_{a_{1},a_{2}}$ over $\mathbb{R}^{N}$ such that
@@
The measure $\mu_{a_{1},a_{2}}^{\mathcal{R}(\theta)}$ is clearly not unique, by the symmetry of $J_{a}^{\mathcal{R}(\theta)}(x)$. By [trimeche], there exists a signed measure $\mu^{\mathcal{R}(\theta)}_{a_{1},a_{2}}$ supported over $B(0,\left\lVert a_{1}\right\rVert_{2}+\left\lVert a_{2}\right\rVert_{2})$ such that the equation in ˜1.4 is satisfied. It remains to determine whether this measure is nonnegative; more precisely, we must determine whether the symmetric version of this measure is nonnegative.
The paper [rosler_product] shows that ˜1.4 holds in the context of radially symmetric Bessel functions. Also, note that it has already been determined that the analogous conjecture for the nonsymmetric eigenfunctions of the Dunkl operators rather than the symmetric eigenfunctions is false, see [convolution_maximal].
As corollaries of the main results of this paper, we prove the following results about the weak convergences of the measures mentioned in ˜1.4. For the proofs of the corollaries, see Subsection˜10.1.
@@
Assume that ˜1.4 is true. Furthermore, assume that $\theta\geq 0$ for all $N\geq 1$ and $\lim_{N\rightarrow\infty}\theta N=\infty$. Let $\mu_{a}$ and $\mu_{b}$ be compactly supported distributions over $\mathbb{R}$. Suppose $a(N),b(N)\in\mathbb{R}^{N}$ for $N\geq 1$ such that $\sum_{i=1}^{N}\frac{1}{N}\delta\left(\frac{a(N)_{i}}{\theta N}\right)\rightarrow\mu_{a}$ and $\sum_{i=1}^{N}\frac{1}{N}\delta\left(\frac{b(N)_{i}}{\theta N}\right)\rightarrow\mu_{b}$ in terms of moments as $N\rightarrow\infty$. Let $\mu$ be the free convolution of $\mu_{a}$ and $\mu_{b}$ as defined in [nica_speicher_2006]*Definition 12.1.
@@
    The distribution $\operatorname{\mathbb{E}}_{a\sim\mu_{a(N),b(N)}^{A^{N-1}(\theta)}}\left[\sum_{i=1}^{N}\frac{1}{N}\delta\left(\frac{a_{i}}{\theta N}\right)\right]$ always converges weakly to $\mu$ as $N\rightarrow\infty$.
@@
    The distribution $\operatorname{\mathbb{E}}_{a\sim\mu_{a(N),b(N)}^{D^{N}(\theta)}}\left[\sum_{i=1}^{N}\frac{1}{2N}\delta\left(\frac{a_{i}}{\theta N}\right)+\frac{1}{2N}\delta\left(-\frac{a_{i}}{\theta N}\right)\right]$ always converges weakly to $\tilde{\mu}$ as $N\rightarrow\infty$.
Remark 1.6.
By “always”, we mean that the statement is true for any choices of the measures $\mu_{a(N),b(N)}^{A^{N-1}(\theta)}$ and $\mu_{a(N),b(N)}^{D^{N}(\theta)}$, since there may be multiple choices.
@@
Assume that ˜1.4 is true. Furthermore, assume that $\theta_{0},\theta_{1}\geq 0$ for all $N\geq 1$, $\lim_{N\rightarrow\infty}\theta_{0}N=\infty$, and $\lim_{N\rightarrow\infty}\frac{\theta_{1}}{\theta_{0}N}=c\in\mathbb{C}$. Let $\mu_{a}$ and $\mu_{b}$ be compactly supported distributions over $\mathbb{R}$. Suppose $a(N),b(N)\in\mathbb{R}^{N}$ for $N\geq 1$ such that $\sum_{i=1}^{N}\frac{1}{N}\delta\left(\frac{a(N)_{i}}{\theta_{0}N}\right)\rightarrow\mu_{a}$ and $\sum_{i=1}^{N}\frac{1}{N}\delta\left(\frac{b(N)_{i}}{\theta_{0}N}\right)\rightarrow\mu_{b}$ in terms of moments as $N\rightarrow\infty$.
Furthermore, define $\mu_{a}^{-}$ and $\mu_{b}^{-}$ by $\mu_{a}^{-}(B)\triangleq\mu_{a}(-B)$ and $\mu_{b}^{-}(B)\triangleq\mu_{b}(-B)$, respectively, for all open subsets $B$ of $\mathbb{R}$. Let $\mu$ be the rectangular free convolution with $\lambda$ set as $\frac{1}{1+c}$ of $\frac{1}{2}(\mu_{a}+\mu_{a}^{-})$ and $\frac{1}{2}(\mu_{b}+\mu_{b}^{-})$ as defined in [rectangular_free_convolution]*Proposition-Definition 2.1. The distribution $\operatorname{\mathbb{E}}_{a\sim\mu^{BC^{N}(\theta_{0},\theta_{1})}_{a(N),b(N)}}\left[\sum_{i=1}^{N}\frac{1}{2N}\delta\left(\frac{a_{i}}{\theta_{0}N}\right)+\frac{1}{2N}\delta\left(-\frac{a_{i}}{\theta_{0}N}\right)\right]$ always converges weakly to $\mu$ as $N\rightarrow\infty$.
We also note that analogues of part (A) of Corollary˜1.5 and Corollary˜1.7 have been established in the $\theta N\rightarrow c\in\mathbb{C}$ and $\theta_{0}N\rightarrow c_{0}\in\mathbb{C},\theta_{1}\rightarrow c_{1}\in\mathbb{C}$ regimes, respectively, see [matrix, rectangularmatrix]. We can similarly establish the analogue of part (B) of Corollary˜1.5 in the $\theta N\rightarrow c\in\mathbb{C}$ regime; its statement is almost the same as that of the analogue of part (A) of the corollary.

### v3

@@ 1.1. Main result
Furthermore, Theorem˜1.1 generalizes the results of [limits_prob_measures] and resolves the question posed in the appendix of [matrix]; see Lemma˜1.2 for a more direct connection with the question that is posed. In particular, [limits_prob_measures] proves the implication of (b) from (a) after modifying the condition that $\lim_{N\rightarrow\infty}\frac{c_{\lambda}(N)}{(\theta N)^{\ell(\lambda)}}=0$ to $|c_{\lambda}(N)|=O(|\theta N|\max(N,|\theta N|)^{o_{N}(1)})$ for $\lambda\in\Gamma$ such that $\ell(\lambda)\geq 2$. The main contribution of this paper is the set of equivalence relations in the theorem where some conditions are based on scaling the coefficients of the logarithms of the input functions $F_{N}^{A}$, $F_{N}^{BC}$, and $F_{N}^{D}$ by varying powers of $\theta N$. The idea of scaling the coefficients by varying powers of $N$ while $\theta$ is fixed is mentioned in the appendix of [matrix]. However, in [limits_prob_measures], the coefficients are only scaled by $(\theta N)^{-1}$ and it is similarly the case that $\lim_{N\rightarrow\infty}|\theta N|=\infty$.
@@
We can also restate Theorem˜1.1 in terms of the Taylor series expansions over the standard basis $M_{\epsilon}$ for $\epsilon\in\mathcal{P}_{N}$ and the coefficients $c_{\epsilon\lambda}$ for $\epsilon,\lambda\in\Gamma$. See Subsection˜2.2 for the definitions of these notions.
Lemma 1.2.
Repeat the statement of Theorem˜1.1. Suppose
$F_{N}^{A}(x_{1},\ldots,x_{N})=\exp\left(\sum_{\epsilon\in\mathcal{P}_{N}}m^{A}_{\epsilon}(N)M_{\epsilon})\right)\in\mathbb{C}[[x_{1},\ldots,x_{N}]],$
$F_{N}^{BC}(x_{1},\ldots,x_{N})=\exp\left(\sum_{\epsilon\in\mathcal{P}_{\text{even};\,N}}m^{BC}_{\epsilon}(N)M_{\epsilon}\right)\in\mathbb{C}[[x_{1}^{2},\ldots,x_{N}^{2}]],$
$F_{N}^{D}(x_{1},\ldots,x_{N})=\exp\left(\sum_{\epsilon\in\mathcal{P}_{\text{even};\,N}}m^{D}_{\epsilon}(N)M_{\epsilon}\right)+e\exp\left(\sum_{\epsilon\in\mathcal{P}_{\text{even};\,N}}l^{D}_{\epsilon}(N)M_{\epsilon}\right)$
$\in\mathbb{C}[[x_{1}^{2},\ldots,x_{N}^{2}]]+e\mathbb{C}[[x_{1}^{2},\ldots,x_{N}^{2}]].$
Consider the following statements.
  1. (a’)
    For all $\lambda\in\Gamma$, $\lim_{N\rightarrow\infty}\frac{\sum_{\epsilon\in\mathcal{P}_{N}}c_{\epsilon\lambda}m^{A}_{\epsilon}(N)}{\theta N}=c_{\lambda}\in\mathbb{C}$ if $\ell(\lambda)=1$ and $\lim_{N\rightarrow\infty}\\ \frac{\sum_{\epsilon\in\mathcal{P}_{N}}c_{\epsilon\lambda}m^{A}_{\epsilon}(N)}{(\theta N)^{\ell(\lambda)}}=0$ if $\ell(\lambda)\geq 2$.
  2. (a”)
    For all $\lambda\in\Gamma$, $\lim_{N\rightarrow\infty}\frac{\sum_{\epsilon\in\mathcal{P}_{N}}c_{\epsilon\lambda}(\epsilon!)^{-1}[1]\partial^{\epsilon}F_{N}^{A}}{\theta N}=c_{\lambda}\in\mathbb{C}$ if $\ell(\lambda)=1$ and $\lim_{N\rightarrow\infty}\\ \frac{\sum_{\epsilon\in\mathcal{P}_{N}}c_{\epsilon\lambda}(\epsilon!)^{-1}[1]\partial^{\epsilon}F_{N}^{A}}{(\theta N)^{\ell(\lambda)}}=0$ if $\ell(\lambda)\geq 2$.
  3. (c’)
    For all $\lambda\in\Gamma_{\text{even}}$, $\lim_{N\rightarrow\infty}\frac{\sum_{\epsilon\in\mathcal{P}_{\text{even};\,N}}c_{\epsilon\lambda}m^{BC}_{\epsilon}(N)}{\theta N}=c_{\lambda}\in\mathbb{C}$ if $\ell(\lambda)=1$ and $\lim_{N\rightarrow\infty}\\ \frac{\sum_{\epsilon\in\mathcal{P}_{\text{even};\,N}}c_{\epsilon\lambda}m^{BC}_{\epsilon}(N)}{(\theta N)^{\ell(\lambda)}}=0$ if $\ell(\lambda)\geq 2$.
  4. (c”)
    For all $\lambda\in\Gamma_{\text{even}}$, $\lim_{N\rightarrow\infty}\frac{\sum_{\epsilon\in\mathcal{P}_{\text{even};\,N}}c_{\epsilon\lambda}(\epsilon!)^{-1}[1]\partial^{\epsilon}F_{N}^{BC}}{\theta N}=c_{\lambda}\in\mathbb{C}$ if $\ell(\lambda)=1$ and $\lim_{N\rightarrow\infty}\frac{\sum_{\epsilon\in\mathcal{P}_{\text{even};\,N}}c_{\epsilon\lambda}(\epsilon!)^{-1}[1]\partial^{\epsilon}F_{N}^{BC}}{(\theta N)^{\ell(\lambda)}}=0$ if $\ell(\lambda)\geq 2$.
  5. (e’)
    For all $\lambda\in\Gamma_{\text{even}}$, $\lim_{N\rightarrow\infty}\frac{\sum_{\epsilon\in\mathcal{P}_{\text{even};\,N}}c_{\epsilon\lambda}m^{D}_{\epsilon}(N)}{\theta N}=c_{\lambda}\in\mathbb{C}$ if $\ell(\lambda)=1$ and $\lim_{N\rightarrow\infty}\\ \frac{\sum_{\epsilon\in\mathcal{P}_{\text{even};\,N}}c_{\epsilon\lambda}m^{D}_{\epsilon}(N)}{(\theta N)^{\ell(\lambda)}}=0$ if $\ell(\lambda)\geq 2$.
  6. (e”)
    For all $\lambda\in\Gamma_{\text{even}}$, $\lim_{N\rightarrow\infty}\frac{\sum_{\epsilon\in\mathcal{P}_{\text{even};\,N}}c_{\epsilon\lambda}(\epsilon!)^{-1}[1]\partial^{\epsilon}F_{N}^{D}}{\theta N}=c_{\lambda}\in\mathbb{C}$ if $\ell(\lambda)=1$ and $\lim_{N\rightarrow\infty}\frac{\sum_{\epsilon\in\mathcal{P}_{\text{even};\,N}}c_{\epsilon\lambda}(\epsilon!)^{-1}[1]\partial^{\epsilon}F_{N}^{D}}{(\theta N)^{\ell(\lambda)}}=0$ if $\ell(\lambda)\geq 2$.
  7. (f’)
    For all $\lambda\in\Gamma_{\text{even}}$, $\lim_{N\rightarrow\infty}\frac{\sum_{\epsilon\in\mathcal{P}_{\text{even};\,N}}c_{\epsilon\lambda}l^{D}_{\epsilon}(N)}{\theta N}=d_{\lambda}\in\mathbb{C}$ if $\ell(\lambda)=1$ and $\lim_{N\rightarrow\infty}\\ \frac{\sum_{\epsilon\in\mathcal{P}_{\text{even};\,N}}c_{\epsilon\lambda}l^{D}_{\epsilon}(N)}{(\theta N)^{\ell(\lambda)}}=0$ if $\ell(\lambda)\geq 2$.
  8. (f”)
    For all $\lambda\in\Gamma_{\text{even}}$, $\lim_{N\rightarrow\infty}\frac{\sum_{\epsilon\in\mathcal{P}_{\text{even};\,N}}c_{\epsilon\lambda}((\epsilon+(1,\ldots,1))!)^{-1}[1]\partial^{\epsilon+(1,\ldots,1)}F_{N}^{D}}{\theta N}=d_{\lambda}\in\mathbb{C}$ if $\ell(\lambda)=1$ and $\lim_{N\rightarrow\infty}\frac{\sum_{\epsilon\in\mathcal{P}_{\text{even};\,N}}c_{\epsilon\lambda}((\epsilon+(1,\ldots,1))!)^{-1}[1]\partial^{\epsilon+(1,\ldots,1)}F_{N}^{D}}{(\theta N)^{\ell(\lambda)}}=0$ if $\ell(\lambda)\geq 2$.
Then, (a’) and (a”), (c’) and (c”), (e’) and (e”), and (f’) and (f”) are equivalent to (a), (c), (e), and (f), respectively.
Remark 1.3.
The partial derivative $\partial^{\epsilon}$ for $\epsilon\in\mathcal{P}_{N}$ denotes $\prod_{i=1}^{\ell(\epsilon)}\partial_{i}^{\epsilon_{i}}$. We can clearly replace it with $\prod_{i=1}^{\ell(\epsilon)}\partial_{\pi(i)}^{\epsilon_{i}}$ for any injection $\pi:[\ell(\epsilon)]\rightarrow[N]$ since the functions $F_{N}^{A}$, $F_{N}^{BC}$, and $F_{N}^{D}$ are symmetric. Furthermore, in (f”), $(1,\ldots,1)$ consists of $N$ copies of $1$ and $(\epsilon+(1,\ldots,1))!=\prod_{i=1}^{\ell(\epsilon)}(\epsilon_{i}+1)$.
Proof of Lemma˜1.2.
The equivalences involving (a’), (c’), (e’), and (f’) are straightforward by the definition of the coefficients $c_{\epsilon\lambda}$. The equivalences involving (a”), (c”), (e”), and (f”) follow from $m_{\epsilon}^{A}(N)=(\epsilon!)^{-1}[1]\partial^{\epsilon}F_{N}^{A}$, $m_{\epsilon}^{BC}(N)=(\epsilon!)^{-1}[1]\partial^{\epsilon}F_{N}^{BC}$, $m_{\epsilon}^{D}(N)=(\epsilon!)^{-1}[1]\partial^{\epsilon}F_{N}^{D}$, and $l_{\epsilon}^{D}(N)=((\epsilon+(1,\ldots,1))!)^{-1}[1]\partial^{\epsilon+(1,\ldots,1)}F_{N}^{D}$. ∎
@@ 1.2. The $\theta N\rightarrow c\in\mathbb{C}$ regime
Note that [matrix] considers regime 3 for the root system $A^{N-1}$ while [rectangularmatrix] considers regime 4. The approach that we use to prove Theorem˜1.1 over regimes 1 and 2 is easily adaptable to proving similar results over regimes 3 and 4 as well. Theorems that we prove over regimes 3 and 4 include Theorem˜6.17, which generalizes the results of [matrix], and Theorem˜7.8, which generalizes the results of [rectangularmatrix]. We also prove Theorem˜8.26 for regime 3 and the type D root system.
@@ 1.4. Weak convergence to the free convolution
From [dunkloperators, dunklbessel], if $\theta\in\theta(\mathcal{R})$ such that $\theta(r)$ has nonnegative real part for all $r\in\mathcal{R}$, then $\theta\in\Theta(\mathcal{R})$, or equivalently, $\mathcal{D}(\mathcal{R}(\theta))$ is invertible. See Subsection˜2.1 for the definitions of these notions. The invertibility of the Dunkl operator implies the existence of it associated eigenfunctions, see Theorems˜2.5 and 2.8. In particular, this implies that the eigenfunctions exist in the context of Theorem˜1.4.
@@
Suppose $\geq 1, N\mathcal{R}\subset\mathbb{R}^{N}$ is a finite root system, and $\theta\in\theta(\mathcal{R})$ is nonnegative. Suppose $a_{1},a_{2}\in\mathbb{R}^{N}$. There exists a nonnegative Borel probability measure $\mu^{\mathcal{R}(\theta)}_{a_{1},a_{2}}$ over $\mathbb{R}^{N}$ such that
@@
The measure $\mu_{a_{1},a_{2}}^{\mathcal{R}(\theta)}$ is clearly not unique, by the symmetry of $J_{a}^{\mathcal{R}(\theta)}(x)$. By [trimeche], there exists a signed measure $\mu^{\mathcal{R}(\theta)}_{a_{1},a_{2}}$ supported over $B(0,\left\lVert a_{1}\right\rVert_{2}+\left\lVert a_{2}\right\rVert_{2})$ such that the equation in Conjecture˜1.6 is satisfied. It remains to determine whether this measure is nonnegative; more precisely, we must determine whether the symmetric version of this measure is nonnegative.
The paper [rosler_product] shows that Conjecture˜1.6 holds in the context of radially symmetric Bessel functions. Also, note that it has already been determined that the analogous conjecture for the nonsymmetric eigenfunctions of the Dunkl operators rather than the symmetric eigenfunctions is false, see [convolution_maximal].
As corollaries of the main results of this paper, we prove the following results about the weak convergences of the measures mentioned in Conjecture˜1.6. For the proofs of the corollaries, see Subsection˜10.1.
@@
Assume that $\theta\geq 0$ for all $N\geq 1$ and $\lim_{N\rightarrow\infty}\theta N=\infty$. Let $\mu_{a}$ and $\mu_{b}$ be compactly supported Borel probability measures over $\mathbb{R}$. Suppose $a(N),b(N)\in\mathbb{R}^{N}$ for $N\geq 1$ such that $\sum_{i=1}^{N}\frac{1}{N}\delta\left(\frac{a(N)_{i}}{\theta N}\right)\rightarrow\mu_{a}$ and $\sum_{i=1}^{N}\frac{1}{N}\delta\left(\frac{b(N)_{i}}{\theta N}\right)\rightarrow\mu_{b}$ in terms of moments as $N\rightarrow\infty$. Let $\mu$ be the free convolution of $\mu_{a}$ and $\mu_{b}$ as defined in [nica_speicher_2006]*Definition 12.1.
@@
    Assume that for all $N\geq 1$, the measure $\mu_{a(N),\,b(N)}^{A^{N-1}(\theta)}$ satisfies the conditions of Conjecture˜1.6. Then, $\operatorname{\mathbb{E}}_{a\sim\mu_{a(N),\,b(N)}^{A^{N-1}(\theta)}}\left[\sum_{i=1}^{N}\frac{1}{N}\delta\left(\frac{a_{i}}{\theta N}\right)\right]$ converges to $\mu$ as $N\rightarrow\infty$ in distribution and in terms of moments.
@@
    Assume that for all $N\geq 1$, the measure $\mu_{a(N),\,b(N)}^{D^{N}(\theta)}$ satisfies the conditions of Conjecture˜1.6. Then, $\operatorname{\mathbb{E}}_{a\sim\mu_{a(N),\,b(N)}^{D^{N}(\theta)}}\left[\sum_{i=1}^{N}\frac{1}{2N}\delta\left(\frac{a_{i}}{\theta N}\right)+\frac{1}{2N}\delta\left(-\frac{a_{i}}{\theta N}\right)\right]$ converges to $\tilde{\mu}$ as $N\rightarrow\infty$ in distribution and in terms of moments.
Conjecture˜1.6 is known for $\theta\in\{\frac{1}{2},1,2\}$. By [crystalrandommat]*Proposition 2.3, the measure $\mu_{a(N),\,b(N)}^{A^{N-1}(\theta)}$ is given by the spectrum of the addition of two random matrices, one with spectrum given by $a(N)$ and the other with spectrum given by $b(N)$. When $\theta\in\{\frac{1}{2},1,2\}$, the matrix is uniformly random and self-adjoint with real, complex, and quaternion entries, respectively. Then, by [voiculescu], see also [matrix]*Theorem 1.1, we can deduce that the measure weakly converges to the free convolution.
@@
Assume that $\theta_{0},\theta_{1}\geq 0$ for all $N\geq 1$, $\lim_{N\rightarrow\infty}\theta_{0}N=\infty$, and $\lim_{N\rightarrow\infty}\\ \frac{\theta_{1}}{\theta_{0}N}=c\in\mathbb{C}$. Let $\mu_{a}$ and $\mu_{b}$ be compactly supported Borel probability measures over $\mathbb{R}$. Suppose $a(N),b(N)\in\mathbb{R}^{N}$ for $N\geq 1$ such that $\sum_{i=1}^{N}\frac{1}{N}\delta\left(\frac{a(N)_{i}}{\theta_{0}N}\right)\rightarrow\mu_{a}$ and $\sum_{i=1}^{N}\frac{1}{N}\\ \delta\left(\frac{b(N)_{i}}{\theta_{0}N}\right)\rightarrow\mu_{b}$ in terms of moments as $N\rightarrow\infty$.
Furthermore, define $\mu_{a}^{-}$ and $\mu_{b}^{-}$ by $\mu_{a}^{-}(B)\triangleq\mu_{a}(-B)$ and $\mu_{b}^{-}(B)\triangleq\mu_{b}(-B)$, respectively, for all open subsets $B$ of $\mathbb{R}$. Let $\mu$ be the rectangular free convolution with $\lambda$ set as $\frac{1}{1+c}$ of $\frac{1}{2}(\mu_{a}+\mu_{a}^{-})$ and $\frac{1}{2}(\mu_{b}+\mu_{b}^{-})$ as defined in [rectangular_free_convolution]*Proposition-Definition 2.1.
Assume that for all $N\geq 1$, the measure $\mu_{a(N),\,b(N)}^{BC^{N}(\theta_{0},\theta_{1})}$ satisfies the conditions of Conjecture˜1.6. Then, $\operatorname{\mathbb{E}}_{a\sim\mu^{BC^{N}(\theta_{0},\theta_{1})}_{a(N),\,b(N)}}\left[\sum_{i=1}^{N}\frac{1}{2N}\delta\left(\frac{a_{i}}{\theta_{0}N}\right)+\frac{1}{2N}\delta\left(-\frac{a_{i}}{\theta_{0}N}\right)\right]$ converges to $\mu$ as $N\rightarrow\infty$ in distribution and in terms of moments.
We also note that analogues of part (A) of Corollary˜1.7 and Corollary˜1.8 have been established in the $\theta N\rightarrow c\in\mathbb{C}$ and $\theta_{0}N\rightarrow c_{0}\in\mathbb{C},\theta_{1}\rightarrow c_{1}\in\mathbb{C}$ regimes, respectively, see [matrix, rectangularmatrix]. We can similarly establish the analogue of part (B) of Corollary˜1.7 in the $\theta N\rightarrow c\in\mathbb{C}$ regime; its statement is almost the same as that of the analogue of part (A) of the corollary. As expected, we can recover the free convolution from the convolution for the $\theta N\rightarrow c\in\mathbb{C}$ regime by taking $|c|\rightarrow\infty$, see [matrix]*Theorem 8.7.
Furthermore, we discuss generalizations of Corollaries˜1.7 and 1.8 in Subsection˜10.5 when $a(N)$ and $b(N)$ are randomly distributed based on exponentially decaying measures rather than deterministic. In particular, we use this framework to analyze the Dyson Brownian motion with random initial values in Example˜10.21. We establish that if the average of the scaled empirical distribution of the initial values converge in terms of moments to a Borel probability measure $\eta$ over $\mathbb{R}$, then the average of the scaled empirical distribution after $\alpha N$ time steps converges in terms of moments to a measure determined by $\eta$ and $\alpha$.